OpenStudy (anonymous):

In a certain county, the number of charter schools is 10 less than twice the number of alternative schools. We know that there are 46 charter schools in the county. How many alternative schools are in the county? 28 56 18 40

OpenStudy (anonymous):

28. you need explanation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

they really struggle to come up with these "word problems" don't they?

OpenStudy (anonymous):

so you work backwards. you take 46 and add 10 because it says "10 less than"

OpenStudy (anonymous):

tell me about it

OpenStudy (anonymous):

you get 56 .now since it says "twice the number" you do the opposite and divide 56 by 2. and voila

OpenStudy (nurali):

Let c = the number of charter schools. Let a = the number of alternative schools. c = 2*a - 10. Number of charter schools is 46. This makes our equation equal to: 46 = 2*a - 10 Add 10 to both sides of this equation to get: 56 = 2*a Divide both sides of this equation by 2 to get: a = 28 There are 28 alternative schools in the country.

OpenStudy (anonymous):

or is it viola?

OpenStudy (anonymous):

viola is an instrument

OpenStudy (anonymous):

right lol.